In a random sample of 16 senior-level chemical engineers, the mean annual earnings was 120250 and the standard deviation was 34700. Assume the annual earnings are normally distributed and construct a 95% confidence interval for the population mean annual earnings for senior-level chemical engineers.

1. The critical value:

2. The standard error of the sample mean:

3. The margin of error:

4. The lower limit of the interval:


5. The upper limit of the interval:
Answer #1

Solution :

Given that,

Point estimate = sample mean = = 120250

sample standard deviation = s = 34700

sample size = n = 16

Degrees of freedom = df = n - 1 = 15

1)

At 95% confidence level the t is ,

= 1 - 95% = 1 - 0.95 = 0.05

/ 2 = 0.05 / 2 = 0.025

t /2,df = t0.025,15 = 2.131

2)

The standard error of the sample mean:

s = (s /n)

= 34700 / 16

= 8675

Margin of error = E = t/2,df * s

= 2.131 * 8675

= 18486.425

The 95% confidence interval estimate of the population mean is,

- E < < + E

120250 - 18486.425 < < 120250 + 18486.425

101763.575 < < 138736.425

3)

The lower limit of the interval: 101763.575

4)

The upper limit of the interval: 138736.425